OpenStudy (anonymous):

multiply out (.5a + .5b)(.5a + .5b)

OpenStudy (anonymous):

\[\Large (5a+5b)^2=(5a)^2+2\cdot(5a\cdot 5b)+(5b)^2\]

OpenStudy (anonymous):

\[\Large (25a^2+50ab+25b^2)\]